% dissociation of 0.024M solution of weak acid HA
(ka = 2 × 10-3) is :
 

Grade:11

1 Answers

Vikas TU
7 years ago
                 HA ============>         H+    +    A-
t=0          0.024                                  0            0
t=t           0.024(1-a)                       0.024a        0.024a
Now from ionic eq. equal conc. method,
[H+] = [A-]
[H+]^2 = Ka*[HA] => 0.024*2*10^-3 => 48 x 10^-6
[H+] = root48) * 10^-3 M
 
Now,
Ca = root(48)*10^-3
a = root(48)/24 => 0.288*100 => 28.88%